Using vertical analysis, prepare a common size comparative balance sheet. (Round percentages to 0 decimal places, eg, 12%.) Delta COMPANY Comparative Balance Sheet December 31, 2017 2016 Amount Percentage Amounft $280 520 ssets Percentage Current assets $450 Plant assets 550 Total assets Liabilities and stockholders equity Current liabilities Long-term debt Common stock $1,000 $800 $120 160 320 200 $180 250 310 Retained earnings 260 Total liabilities and stockholders equity $1,000 $800
Answer #1

Under vertical analysis;each item will be represented as a % of total assets/Total liabilities and stockholders equity.

2017 2016
Amount % Amount %
Assets
Current Assets 450 (450/1000)=45% 280 (280/800)=35%
Plant Assets 550 (550/1000)=55% 520 (520/800)=65%
Total Assets $1000 100% $800 100%
Liabilities and Stockholders equity
Current Liabilities 180 (180/1000)=18% 120 (120/800)=15%
Long-term debt 250 (250/1000)=25% 160 (160/800)=20%
Common Stock 310 (310/1000)=31% 320 (320/800)=40%
Retained earnings 260 (260/1000)=26% 200 (200/800)=25%
Total Liabilities and Stockholders equity 1000 100% 800 100%